OpenStudy (sissyedgar):

What is the slope-intercept form of the function that contains the points (3, 1) and (2, 9)? y = ___ x + ___ Thankiz!!!! ღ

OpenStudy (sissyedgar):

@confluxepic

OpenStudy (astrophysics):

They want your answer in y = mx+b form you can get the slope simply by using \[ m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] and you can set it up in point - slope form \[y-y _{1}=m(x-x _{1})\] where your coordinates for x1 and y1 are the first point (3,1) do some little algebra you'll end up with slope intercept form.

OpenStudy (sissyedgar):

@TheSmartOne Im so lost rn! xP

OpenStudy (astrophysics):

Naw you just didn't want to read and do it yourself

OpenStudy (sissyedgar):

Nurp. im just lost

TheSmartOne (thesmartone):

Whatever help I can give will basically be similar to what my buddy Astro said :P